# Mathematics with large language models as provers and verifiers

### v2

@@ Abstract
During 2024 and 2025 the discussion about the theorem-proving capabilities of large language models started reporting interesting success stories, mostly to do with difficult exercises (such as problems from the International Mathematical Olympiad), but also with conjectures [8] formulated for the purpose of verifying whether the artificial intelligence could prove it. In this paper we report a theorem proving feat achieved by ChatGPT by using a protocol involving different prover and verifier instances of the gpt-5 model working collaboratively. To make sure that the produced proofs do not suffer from hallucinations, the final proof is formally verified by the lean proof assistant, and the conformance of premises and conclusion of the lean code is verified by a human. Our methodology is by no means complete or exact. It was nonetheless able to solve five out of six 2025 IMO problems, and close about a third of the sixty-six number theory conjectures in [4].
@@ 1 Introduction
This paper describes a protocol, involving the OpenAI Application Programming Interface (API) with a minimal human interaction, which was able to solve five out of six problems of the International Mathematical Olympiads (IMO), close twenty-two out of sixty-six open conjectures on cyclic numbers proposed in [4], as well as discover and prove several new theorems in a variety of fields of mathematics. The human intervention only arises to verify that a formal version of the generated theorem statement is consistent with its semi-formal natural language description. More precisely, we only ask the human to verify the conformance of all premises and the conclusion of the formal and the semi-formal versions of the proof [^1] At the time of publishing this technical report we are still carrying out the human part of the protocol: many proofs should not be trusted yet; but those we reported in this paper appear correct to us (as human verifiers)..
@@ 1.1 Contributions of this paper
The main contribution of this paper is a conceptually simple protocol, involving the repeated interaction of different LLM-based computational agents and terminating with a final human verification, that proves theorems. This protocol solved five out of the six problems of the 2025 IMO, exactly like the specialized (and as far as we know undisclosed) LLMs built by Google DeepMind and OpenAI; it solved a third of the number theoretical conjectures in [4]; and it discovered and settled new conjectures in three fields of mathematics (see Sect. 4.1.1 and 5.3). Obviously, the protocol does not guarantee that every conjecture will be proved or disproved: after a given amount of time it can give up empty-handed.
@@
A more subtle contribution, which may not sound new to LLM experts but may surprise traditional computer scientists, is the idea that “programming with an LLM” involves rhetorical skills in giving natural language instructions to LLM-based agents to tell them how to behave. Much like the grammarians of the late middle ages who fought over the vocative of ego, LLM programmers often resort to various imperative forms “to be”: we tell our LLM agents “you are a meticulous researcher”, “be cautious”, “be concise”, and even “you are a senior research assistant who is finding advanced sources in the literature to tackle an open problem”. These prompts are passed to LLM instances to prepare them for their work and define their output (see Sect. 4.1). The new field of computational pragmatics is arising from such needs [2, 22, 17].
@@ 2 A review of the relevant literature
Consider this scenario: the LLM generates a proof and the corresponding proof assistant code, both of which may be subject to hallucinations. The hallucinated code may still be accepted as valid by the proof assistant as long as the premises lead to the conclusion by inference, but these premises may not correspond to those of the LLM-generated proof. While we agree with [19, 33] that such an occurrence is unlikely, we also think that a wrong proof of this kind would be extremely difficult to catch by humans exactly because it is known to have been checked by a proof assistant. Our process overcomes this difficulty by introducing a very limited human interaction: a human verifies the conformance of all premises and conclusion of the proof between the proof assistant code and the LLM-generated proof [^2] Humans also make mistakes when proving theorems, but at this stage of LLM technology development we should be happy if computers do not make more mistakes than we do..
The interaction of humans and AI agents in the context of proving theorems is not new. See for example the end of [25, §2] for a few cases. In [25], the interaction between the human and the AI agent is based on the agent being a “co-pilot”: the human leads the investigation but can ask the AI agents for suggestions. In our case the human is limited to a verification task to be carried out only once, at the end of the computer-based process.
@@ 4 System implementation
The protocol presented in Sect. 3 is similar, albeit not identical, to several other approaches in the literature (see Sect. 2). The only theoretically interesting innovation is the minimal interaction with a human right at the end (see Prop. 3.1 and the remarks below it), but this just ensures a correctness guarantee: it cannot explain the success of our protocol in solving open conjectures and discovering new theorems (see Sect. 5).
@@ 4.4 Human interaction issues
In practice, the human interaction remains very time-consuming. This is mostly due to the limited ability shown by gpt-5 in converting its own natural language proofs into formal lean codes. So, although our protocol does minimize human interaction, this interaction is remains excessive, because, aside from a handful of cases, the human has to construct the whole lean code almost from scratch. A more intensive LLM training using lean codes could improve the situation. This is another situation (as that described in Sect. 4.2) which would be better tackled by OpenAI. For the time being, we instructed OpenAI’s Codex interface to reduce to “axiom” any proof step that Codex can find by means of a web search. This reduces the extent of lean checking but makes it actually possible to partly fill the certification part of our protocol.
@@ 5 Results
In this section we give an account of the achievements of our system. We note that gpt-5 was trained on data up to 2024, while the problems we submitted to our system all date from 2025. Therefore, the LLM in our TTVR loop cannot have been contaminated in the sense of [26], namely that its training set cannot contain any IMO 2025 information, nor any solution to conjectures in [4] (the situation for the research mode conjectures in Sect. 5.3 is more fluid in this respect: some contamination could be possible). We also emphasize the fact that our prompts do not include instructions to the LLM to the effect of checking the literature; on the contrary, they discourage the production of bullet lists with literature references.
@@ 5.1.1 Notes about the proofs
    The two proofs are different, even though the proof lengths are similar. The human proof is more compact and geometrically clearer to a human reader. The computer proof is algebraic (instead of geometric) insofar as it uses the complex plane to reason about plane geometry (this was due to an explicit instruction to the LLM on our part through a prompt). It relates to the human proof similarly to the way Miles Edwards’ proof of Heron’s formula relates to most other proofs of the same result (see [18, §2] for more details).
@@ 5.2 Conjectures of primes and cyclics
Our system produced a proof or refutation for twenty-two out of the sixty-six conjectures: namely 3, 6, 9, 14, 17, 20, 32, 35, 36, 37, 41, 42, 47, 50, 51, 52, 53, 54, 56, 59, 60, 61, with seventeen proofs and five refutations (35, 50, 51, 53, 59). Seven out of the sixty-six conjectures, namely 2, 3, 4, 36, 37, 52, 53, had already been closed by C. Pomerance prior to the publication of [4], by personal communication to its author (as far as we know, Pomerance’s proofs remain unpublished to date). The five conjectures solved by Pomerance that appear in our list (3, 36, 37, 52, 53) must therefore be removed from the list of “open conjectures”. This leaves our system with 17 (previously open) conjectures for which it obtained a proof.
conjecture 3 6 9 14 17 20 32 35 36 37 41 42 47 50 51 52 53 54 56 59 60 61
||itn|| 4 9 9 6 1 9 3 3 4 8 15 1 9 1 1 9 1 9 9 1 ? ?
O or C C O O O O O O O C C O O O O O C C O O O O O
P or R P P P P P P P R P P P P P R R P R P P R P P
@@
In Table [^6] Missing entries correspond to ongoing tests. 3 we report, for each proof produced by our system: the number of TTVR iterations ($|\mbox{itn}|$), whether the conjecture was open or already closed (O/C), whether the conjecture was settled by proof or refutation (P/R), whether the authors of this paper (who are not specialists in analytic number theory) think that the semi-formal natural language proofs are correct or not, and whether the lean certificate is correct and conformant with the premises and conclusions of the corresponding proofs.
We note that some of the confirmed conjectures have been “closed” only in the asymptotic sense, not in the absolute one. For example, the proof of Conjecture 6 shows that there is an $n_{0}$ such that for every $n>n_{0}$ the conjecture holds, but we have no idea about this $n_{0}$, so Conjecture 6 is still open “for all $n\in\mathbb{N}$”. This is also the case for Conj. 32.
@@ 5.3 Research mode results
In the rest of this section, output generated from the system (possibly only slightly edited for typos or to merge multiple definitions) is on light gray background.
@@ 6 Conclusion
We described a system that employs the OpenAI LLM gpt-5 in order to prove mathematical theorems. The system is based on protocol including an alternating loop of prover and verifier agents impersonated by gpt-5, followed by the creation of a formal lean code version of the proof (if any was found). A human verifies the conformance of the natural language version with the formal version of the premises and conclusion of the theorem, which ensures proof correctness. We showcase our system’s capabilities: (i) on the International Mathematical Olympiad problems in 2025, where our system was as successful as OpenAI and Google DeepMind’s internal LLMs, (ii) on the solution of about a third out of the 66 conjectures on cyclics proposed in [4], and (iii) on the solution of a list on conjectures, formulated by the system itself, in three fields of mathematics: gradient descent in machine learning, the 1-2-3 conjecture in graph theory, and sphere packings in Euclidean spaces.

### v3

@@ Abstract
During 2024 and 2025 the discussion about the theorem-proving capabilities of large language models started reporting interesting success stories, mostly to do with difficult exercises (such as problems from the International Mathematical Olympiad), but also with conjectures [9] formulated for the purpose of verifying whether the artificial intelligence could prove it. In this paper we report a theorem proving feat achieved by ChatGPT by using a protocol involving different prover and verifier instances of the gpt-5 model working collaboratively. To make sure that the produced proofs do not suffer from hallucinations, the final proof is formally verified by the lean proof assistant, and the conformance of premises and conclusion of the lean code is verified by a human. Our methodology is by no means complete or exact. It was nonetheless able to solve five out of six 2025 IMO problems, and close almost a third of the sixty-six number theory conjectures in [5].
@@ 1 Introduction
This paper describes a protocol, involving the OpenAI Application Programming Interface (API) with a minimal human interaction, which was able to solve five out of six problems of the International Mathematical Olympiads (IMO), close almost a third of sixty-six open conjectures on cyclic numbers proposed in [5], as well as discover and prove several new theorems in a variety of fields of mathematics. The human intervention only arises to verify that a formal version of the generated theorem statement is consistent with its semi-formal natural language description. More precisely, we only ask the human to verify the conformance of all premises and the conclusion of the formal and the semi-formal versions of the proof [^1] At the time of publishing this technical report we are still carrying out the human part of the protocol: no proof should not be trusted yet..
@@ 1.1 Contributions of this paper
The main contribution of this paper is a conceptually simple protocol, involving the repeated interaction of different LLM-based computational agents and terminating with a final human verification, that proves theorems. This protocol solved five out of the six problems of the 2025 IMO, exactly like the specialized (and as far as we know undisclosed) LLMs built by Google DeepMind and OpenAI; it solved several theoretical conjectures in [5]; and it discovered and solved new conjectures in three fields of mathematics (see Sect. 4.1.1 and 5.3) [^1] None of these announced results is sure yet: the proofs have been produced in semi-formal natural language, but the verification process involves a mixture of human and automatic work, which takes longer than initially anticipated. Obviously, the protocol does not guarantee that every conjecture will be proved or disproved: after a given amount of time it can give up empty-handed.
@@
A more subtle contribution, which may not sound new to LLM experts but may surprise traditional computer scientists, is the idea that “programming with an LLM” involves rhetorical skills in giving natural language instructions to LLM-based agents to tell them how to behave. Much like the grammarians of the late middle ages who fought over the vocative of ego, LLM programmers often resort to various imperative forms “to be”: we tell our LLM agents “you are a meticulous researcher”, “be cautious”, “be concise”, and even “you are a senior research assistant who is finding advanced sources in the literature to tackle an open problem”. These prompts are passed to LLM instances to prepare them for their work and define their output (see Sect. 4.1). The new field of computational pragmatics is arising from such needs [3, 25, 18], but most of the literature takes “prompt engineering” to mean different data structures (sequences, trees, tables) containing inter-dependent prompts and corresponding answers, all of which are used to provide additional training to the LLM at hand. The prompts we use in our paper are of the “zero-shot” type with no additional training [19, 1].
@@ 2 A review of the relevant literature
Consider this scenario: the LLM generates a proof and the corresponding proof assistant code, both of which may be subject to hallucinations. The hallucinated code may still be accepted as valid by the proof assistant as long as the premises lead to the conclusion by inference, but these premises may not correspond to those of the LLM-generated proof. While we agree with [21, 36] that such an occurrence is unlikely, we also think that a wrong proof of this kind would be extremely difficult to catch by humans exactly because it is known to have been checked by a proof assistant. Our process overcomes this difficulty by introducing a limited human interaction: a human verifies the conformance of all premises and conclusion of the proof between the proof assistant code and the LLM-generated proof [^3] Humans also make mistakes when proving theorems, but at this stage of LLM technology development we should be happy if computers do not make more mistakes than we do..
The interaction of humans and AI agents in the context of proving theorems is not new. See for example the end of [28, §2] for a few cases. In [28], the interaction between the human and the AI agent is based on the agent being a “co-pilot”: the human leads the investigation but can ask the AI agents for suggestions. In our case the human is limited to a verification task to be carried out only once, at the end of the computer-based process.
@@ 4 System implementation
The protocol presented in Sect. 3 is similar, albeit not identical, to several other approaches in the literature (see Sect. 2). The only theoretically interesting innovation is the limited interaction with a human at the very end (see Prop. 3.1 and the remarks below it), but this just ensures a correctness guarantee: it cannot explain the capability of our protocol for solving open conjectures and discovering new theorem statements (see Sect. 5).
@@ 4.4 Human interaction issues
In practice, the human interaction, albeit limited to a verification at the end, remains very time-consuming. This is mostly due to the limited ability shown by gpt-5 in converting its own natural language proofs into formal lean codes. So, although our protocol does minimize human interaction, this interaction remains considerable, because the human has to construct the whole lean code from a relatively thin base. A more intensive LLM training using lean codes could improve the situation. This is another situation (as that described in Sect. 4.2) which would be better tackled by OpenAI.
In order to reduce human interaction, we had initially instructed OpenAI’s Codex interface to reduce to “axiom” any proof step that it could find by means of a web search. This reduced the extent of lean checking (the formalization was partial) but made it possible to run a limited certification step at the end of our protocol [^6] We have since enlisted the help of human experts in analytic number theory to help us verify the solutions of conjectures in [5]. Some of them were found to be incorrect despite the limited formalization effort; and not all of them have been yet vetted..
@@ 5 Results
In this section we give an account of our system’s achievements. We note that gpt-5 was trained on data up to 2024, while the problems we submitted to our system all date from 2025. Therefore, the LLM in our TTVR loop cannot have been contaminated in the sense of [29], namely that its training set cannot contain any IMO 2025 information, nor any solution to conjectures in [5] (the situation for the research mode conjectures in Sect. 5.3 is more fluid in this respect: some contamination could be possible). We also emphasize the fact that our prompts do not include instructions to the LLM to the effect of checking the literature; on the contrary, they discourage the production of bullet lists with literature references.
@@ 5.1.1 Notes about the proofs
    The two proofs are different, even though the proof lengths are similar. The human proof is more compact and geometrically clearer to a human reader. The computer proof is algebraic (instead of geometric) insofar as it uses the complex plane to reason about plane geometry (this was due to an explicit instruction to the LLM on our part through a prompt). It relates to the human proof similarly to the way Miles Edwards’ proof of Heron’s formula relates to most other proofs of the same result (see [20, §2] for more details).
@@ 5.2 Conjectures of primes and cyclics
Our system produced a proof [^8] In analytic number theory, a “proof” can only be correct, otherwise it is not a proof, but a fallacious argument. An “argument” may be correct or incorrect, and a correct argument provides a proof. Moreover, one does not use “theorem” to just indicate the theorem statement: a theorem is the statement followed by its proof. Because the subject matter of this paper is the automated generation of proofs rather than analytic number theory, we abuse those rules to keep descriptions more compact: proofs and refutations may be correct or incorrect. or refutation for twenty out of the sixty-six conjectures: namely 3, 6, 9, 14, 17, 20, 32, 35, 36, 37, 41, 42, 47, 52, 53, 54, 56, 59, 60, 61, with seventeen proofs and three refutations (35, 53, 59). Seven out of the sixty-six conjectures, namely 2, 3, 4, 36, 37, 52, 53, had already been closed by C. Pomerance [23] prior to the publication of [5], by personal communication to its author. The five conjectures solved by Pomerance that appear in our list (3, 36, 37, 52, 53) must therefore be removed from the list of “open conjectures”. This leaves our system with 17 (previously open) conjectures for which it obtained a proof.
conjecture 3 6 9 14 17 20 32 35 36 37 41 42 47 52 53 54 56 59 60 61
||itn|| 4 9 9 6 1 9 3 3 4 8 15 1 9 9 1 9 9 1 1 1
O or C C O O O O O O O C C O O O C C O O O O O
P or R P P P P P P P R P P P P P P R P P R P P
@@
In Table 3 we report, for each proof produced [^9] Missing entries in Table 3 correspond to ongoing tests. by our system: the number of TTVR iterations ($|\mbox{itn}|$), whether the conjecture was open or already closed (O/C), whether the conjecture was settled by proof or refutation (P/R), whether the authors of this paper (who are not specialists in analytic number theory [^10] These proofs ar enow being vetted by experts in the subject matters, who have already invalidated some of the proofs we believed to be correct, and have yet to go through the complete list.) think that the semi-formal natural language proofs are correct or not, and whether the lean certificate is correct and conformant with the premises and conclusions of the corresponding proofs.
We note that some of the confirmed conjectures have been “closed” only partially, i.e. in the asymptotic sense, not in the absolute one. For example, the proof of Conjecture 6 shows that there is an $n_{0}$ such that for every $n>n_{0}$ the conjecture holds, but we have no idea about this $n_{0}$, so Conjecture 6 is still open “for all $n\in\mathbb{N}$”. This is also the case for Conj. 32.
@@ 5.3 Research mode results
In the rest of this section, output generated from the system (possibly only slightly edited for typos or to merge multiple definitions) is on light gray background [^12] None of the proofs in this section has yet been verified formally in lean, nor has it been verified by human experts in the subject matters of the three topics..
@@ 6 Conclusion
We described a system that employs the OpenAI LLM gpt-5 in order to prove mathematical theorems. The system is based on protocol including an alternating loop of prover and verifier agents impersonated by gpt-5, followed by the creation of a formal lean code version of the proof (if any was found). A human verifies the conformance of the natural language version with the formal version of the premises and conclusion of the theorem, which ensures proof correctness. We showcase our system’s capabilities: (i) on the International Mathematical Olympiad problems in 2025, where our system was as successful as OpenAI and Google DeepMind’s internal LLMs, (ii) on the solution of several conjectures on cyclics proposed in [5], and (iii) on the solution of a list on conjectures, formulated by the system itself, in three fields of mathematics: gradient descent in machine learning, the 1-2-3 conjecture in graph theory, and sphere packings in Euclidean spaces.